Please answer back with a detailed description and logical thought process. Please go over 1 through 4

Directions: Solve the following problems and answer the following questions. Then record your responses in the corresponding online test feature, paying careful attention to precise and formatting as instructed or noted For Problems #1 through #8, recall the following context Dr. Sue-Pearl Tennant, is the chief administrator for the Pavonia Point Regional School District, serving students grades P through 8. She is reviewing some presorted Human Resource data about the truncated ages of her teaching staff in district. The following census data documents the ages of all her full-time faculty, rounded down to the nearest whole year 23 32 37 36 38 45 49 58 36 39 46 36 40 46 52 60 32 32 32 36 36 36 37 38 47 57 62 50 59 56 56 58 60 60 61 62 67 67 68 68 72 72 75 76 80 82 85 102 #1 : Compute the following measure of position or location: what data value approximately corresponds to the 15th percentile? [COMMENTS& HINTS Report your answer rounded to the nearest whole number as necessary, such as 1 or 99 or 200. The data value may or may not be an actual data value in this collection.] #2: Compute the following measure of position or location: what approximate percentile corresponds to the data value, 63? [COMMENTS & HINTS: Report your answer rounded to the nearest whole percentile, but expressed as an ordinal number without any percent sign; for example, report 1st or 11th or 12th or 22nd or 43rd or 99th. Do not worry about superscripting the suffixes, but yes, include the proper suffixes. The data value may or may not be an actual data value in this collection.] #3: Compute the following measure of position or location: what data value approximately corresponds to the 80th percentile? [COMMENTS & HINTs Report your answer rounded to the nearest whole number as necessary, such as 1 or 99 or 200. The data value may or may not be an actual data value in this collection.] #4: Compute the following measure of position or location: what approximate percentile corresponds to the data value, 34? [COMMENTS & HINTS: Report your answer rounded to the nearest whole percentile, but expressed as an ordinal number without any percent sign; for example, report 1st or 11th or 12th or 22nd or 43rd or 99th. Do not worry about superscripting the suffixes, but yes, include the proper suffixes. The data value may or may not be an actual data value in this collection.] #5: Compute the following measure of position or location: what data value approximately corresponds to the 44th percentile? [COMMENTS & HINTS Report your answer rounded to the nearest whole number as necessary, such as 1 or 99 or 200 The data value may or may not he an actual data valuue in this
Answer #1

1) start by multiplying 15% times the total number , which gives 15% ? 80 = 0.15*80 = 12(the index).

Counting from the smallest to the largest value in the data set, you go until you find the12th value in the data set. That value is 24, and it’s the 15th percentile for this data set.

2) Here data value is 63, approximate data value in given data set is 62 and corresponding index of this data value are 58 and 59. So index number is = (58+59)/2 = 58.5.

start divide by 58.5 by total number of data 80= 58.5/80 = 0.732 which gives 73% i.e 73 th percentile. Here approximate 73th percentile corresponding to data value 63.

3)

start by multiplying 80% times the total number , which gives 80% ? 80 = 0.80*80 = 64(the index).

Counting from the smallest to the largest value in the data set, you go until you find the 64th value in the data set. That value is 66, and it’s the 80th percentile for this data set.

4)

Here data value is 34, approximate data value in given data set is 32 and corresponding index of this data value are 17,18 and 19. So index number is = (17+18+19)/3 = 18.

start divide by 18 by total number of data 80= 18/80 = 0.225 which approximate 0.23 which gives 23% i.e 23 th percentile.

Here approximate 23th percentile corresponding to data value 34.

5) start by multiplying 44% times the total number , which gives 44% ? 80 = 0.44*80 = 35.2(the index). Rounding up to the nearest whole number, you get 35.

Counting from the smallest to the largest value in the data set, you go until you find the 35th value in the data set. That value is 44, and it’s the 44th percentile for this data set.